# Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

### v3

@@ 4.4 How Does the LoRA Scheme Effectively Fine-tune LLMs?
Given the substantial computational resources required to train LLMs and their powerful generalization abilities, we believe that more efficient use of pre-trained models is essential. Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient. A prominent solution to this challenge is the LoRA (Hu et al. 2021), which can be expressed as follows:
@@
$\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
From Eq. 17, it can be seen that LoRA essentially fine-tunes the UAT parameters layer by layer for a specific task.
@@
4.5 Rethinking LLMs
The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing? Figure 7 illustrates the comparison between the language processing processes of LLMs and humans. Both start with language encoding—humans encode language through a character-based system, while LLMs use numerical arrays. At this level, there is almost no difference. Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using UAT. Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans. From this perspective, the differences between humans and LLMs in language processing seem minimal.
So how do we explain human understanding, analysis of language, and memory retrieval? Are these also capabilities of LLMs? First, we can consider the human brain as a combination of one or more UAT models, which are randomly initialized at birth. What supports such an assumption? The original neural network, the perceptron (Rosenblatt 1963), was designed based on the human neuron. Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons. However, through our derivation, the Transformer can also be understood as an implementation of UAT, and the perceptron can be seen as a layer of UAT. The human brain can be considered as a cluster of multi-layer perception. Learning language in a social environment is akin to training the UATs in the brain. Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs. Understanding and analysis: When we respond to a question posed by someone, it can be seen as a simple example of understanding and analysis, which both humans and LLMs can accomplish. Humans may guess the answer to unfamiliar questions based on their experiences, which can be either correct or incorrect, and LLMs exhibit similar behavior. Memory retrieval: This is essentially fitting specific results based on particular words, as there is no actual database in the brain. For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface. This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E). However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions.
Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face. We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple UAT models to produce reasonable results). Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs. LLMs, on the other hand, are limited to function approximation within the corpora data.
5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s MHA to learn from context.
Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT. By leveraging the theoretical framework provided by the UAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
Appendix A The Property of Diamond
[…]
$\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]$ (1)
$=$ $\mathbf{W}_{1}\diamond[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T}\mathbf{x}=\mathbf{W}_{12}\mathbf{x}$
$=$ $(\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T})^{T}\diamond\mathbf{x}$
$=$ $(\mathbf{W}_{2}\mathbf{W}_{1})\diamond\mathbf{x}$
$=$ $\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
$[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}$ (2)
$=$ $[\mathbf{W}_{1}^{T}\mathbf{x}]\diamond\mathbf{W}_{2}$
$=$ $\mathbf{W}_{2}^{T}[\mathbf{W}_{1}^{T}\mathbf{x}]$
$=$ $[\mathbf{W}_{2}^{T}\mathbf{W}_{1}^{T}]^{T}\diamond\mathbf{x}$
$=$ $[\mathbf{W}_{1}\mathbf{W}_{2}]\diamond\mathbf{x}$
$=$ $\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
@@ Appendix A The Property of Diamond
Figure 2: The process of Transformer.
Appendix B The Matrix-Vector Form of MHA
Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the softmax operation. Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. Figure 1.c shows $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$. In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{HVi}^{\prime}\mathbf{x}_{i}^{\prime}$. This process is divided into four parts: Figure 1.d.1 represents the general form of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1. Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\mathbf{H}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $(\mathbf{W}_{Vi}^{\prime})^{T}\diamond\mathbf{H}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}$ in Figure 1.d.4. Finally, Figure 1.d.5 is obtained as $\mathbf{W}_{HVi}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=\mathbf{H}_{i}^{\prime}$. It can be observed that $\mathbf{W}_{HVi}^{\prime}$ is a dense matrix. In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}$. Figure 1.e.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.e.2 represents the Matrix-vector format of Figure 1.e.1.
Appendix C The UAT Format of Transformer
In this section, we will prove multi-layer Transformer is the implementation of UAT.
C.1 The Properties of UAT
Before expressing Transformers in the UAT format, we present two lemma regarding UAT. There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$. The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$. Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
@@ C.1 The Properties of UAT
$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$ (3)
$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$
Eq. (3) shows the representation of UAT multiplying a matrix. Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the result remains consistent with the original UAT mathematical form. Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
Lemma 2. Let the general term of the network be written as: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, $i=0,1,2,3,\ldots$. The multi-layer network of this form still corresponds to the mathematical form of UAT, and we refer to such a network as a residual network.
To prove that a multi-layer network with the general term written as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ corresponds to the UAT mathematical form, we first provide the forms of single-layer and two-layer networks, as shown in Eq. (4) and (5).
@@
In Eq. (5), let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Eq. (6) can be written as:
@@
Why can we combine the above terms? Once the network training is complete, we know these parameters, so they can be directly calculated and thus combined. And $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$ clearly has the same mathematical form as UAT. We can understand it as dynamically fitting the bias term $\mathbf{b}^{\prime}_{2,2}$ using UAT based on the input. Therefore, we have proven that the mathematical form of one and two-layer residual networks is consistent with the UAT mathematical form.
@@
Assume that the mathematical form of the first $i$ layers of the residual network is consistent with UAT. Our goal is to prove that the mathematical form of the $i+1$-th layer of the residual network is still consistent with UAT. For the convenience of expression, we make the following definition: since the mathematical form of the first $i$ layers of the residual network is consistent with UAT, we write $\mathbf{x}^{\prime}_{i}$ as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$, where the first term is written separately and the rest is written as the remainder term $UAT^{R}_{i}$. Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
@@
First, consider the part $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$. Substitute $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, and we get Eq. (7). Let $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,3}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with UAT, and we have already proven that the mathematical form of UAT remains unchanged when multiplied by a matrix, we have proven that the mathematical form of the first part is consistent with UAT.

### v4

@@ 4.4 How Does the LoRA Scheme Effectively Fine-tune LLMs?
Given the substantial computational resources required to train LLMs and their powerful generalization abilities, we believe that more efficient use of pre-trained models is essential. Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient. A prominent solution to this challenge is the LoRA Hu et al. (2021), which can be expressed as follows:
@@
$+\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
From Eq. 18, it can be seen that LoRA essentially fine-tunes the DUAT parameters layer by layer for a specific task.
5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the DUAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s MHA to learn from context.
Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT. By leveraging the theoretical framework provided by the DUAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
6 Rethinking LLMs
The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing? Figure 7 illustrates the comparison between the language processing processes of LLMs and humans. Both start with language encoding—humans encode language through a character-based system, while LLMs use numerical arrays. At this level, there is almost no difference. Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using DUAT. Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans. From this perspective, the differences between humans and LLMs in language processing seem minimal.
So how do we explain human understanding, analysis of language, and memory retrieval? Are these also capabilities of LLMs? First, we can consider the human brain as a combination of one or more DUAT models, which are randomly initialized at birth. What supports such an assumption? The original neural network, the perceptron Rosenblatt (1963), was designed based on the human neuron. Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons. However, through our derivation, the Transformer can also be understood as an implementation of DUAT, and the perceptron can be seen as a simple UAT. So the human brain can be considered as a cluster of multi-layer perception. Learning language in a social environment is akin to training the DUATs in the brain. Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs. Taking memory retrieval as an example: This is essentially fitting specific results based on particular words, as there is no actual database in the brain. For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface. This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E). However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions (Wang & Li, 2024a).
Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face. We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple DUAT models to produce reasonable results). Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs. LLMs, on the other hand, are limited to function approximation within the corpora data.
@@ 6 Rethinking LLMs
Appendix A The Matrix-Vector Form of MHA
Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the softmax operation. Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. Figure 1.c shows $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$. In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $(\mathbf{W}_{HVi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$. This process is divided into four parts: Figure 1.d.1 represents the general form of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1. Figure 1.d.3 first rewrites Figure 1.d.2 using matrix multiplication as $[\mathbf{H}_{i}^{\prime}\mathbf{W}_{Vi}^{\prime})]^{T}\mathbf{x}_{i}^{\prime}=\hat{\mathbf{H}}_{i}^{\prime}$, and then express it as $(\mathbf{H}_{i}^{\prime})^{T}[\mathbf{W}_{Vi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}]=\hat{\mathbf{H}}_{i}^{\prime}$ in Figure 1.d.4. Finally, Figure 1.d.5 is obtained as $(\mathbf{W}_{HVi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}=(\hat{\mathbf{H}}_{i}^{\prime})^{T}$. It can be observed that $\mathbf{W}_{HVi}^{\prime}$ is a dense matrix. In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=(\hat{\mathbf{H}})^{T}$. Figure 1.e.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.e.2 represents the Matrix-vector format of Figure 1.e.1.
@@ Appendix A The Matrix-Vector Form of MHA
Appendix B Transformer: The implementation of DUAT
In this section, we will prove multi-layer Transformer is the implementation of DUAT.
B.1 The Properties of UAT
Before expressing residual-based CNNs and Transformer-based ViTs in the UAT format, we present a lemma regarding UAT. There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$. The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$. Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
@@ B.1 The Properties of UAT
$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{b}_{j}\right)$ (19)
$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{b}_{j}\right)$
Eq. 19 shows the representation of UAT multiplying a matrix. Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the general mathematical form of Eq. 19 remains consistent with the original UAT mathematical form. Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
B.2 From UAT to DUAT
In this section, we first present a general form for a single-layer residual term in a network and demonstrate that the mathematical form of a multi-layer network composed of this residual term aligns with the DUAT framework. Before proceeding with the formal proof, we first define DUAT explicitly: DUAT shares the overall mathematical form of UAT, but with certain parameters influenced by the input, allowing them to change dynamically in response to it. We refer to these parameters as ”dynamic parameters.” In standard UAT, parameters remain fixed once training is complete, whereas, in DUAT, the dynamic parameters are functions of the input and thus vary with it. Consequently, dynamic parameters in DUAT may take the form of complex functions. Currently, DUAT primarily is implemented by residual structure, meaning that, in general, these complex functions are also DUAT. Next, we will give the proof.
A general residual term of the network can be written as:
$\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,2})$ (20)
where $i=1,2,3,\ldots$ and $\mathbf{x}^{\prime}_{0}$ represents the input. A multi-layer network structured in this way aligns with the mathematical form of DUAT. To demonstrate that a multi-layer network leverages the general term of Eq. 20 corresponds to DUAT, we first examine the forms of single-layer and two-layer networks, as shown in Eqs. 21 and 22.
@@ B.2 From UAT to DUAT
In Eq. 22, let $\mathbf{W}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{2,2}=\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$. Thus, Eq. 22 can be written into Eq. 23.
@@
For the cleatity, we could write Eq. 21 and Eq. 23 into Eq. 24 and Eq. 25, where $UAT^{R}_{1}=\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ and $UAT^{R}_{2}=\Sigma_{j=1}^{2}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$.
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+UAT^{R}_{1}$ (24)
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+UAT^{R}_{2}$ (25)
According to Eq. 24 and Eq. 25, aside from the initial term, the overall mathematical forms of the residual terms $UAT^{R}_{1}$ and $UAT^{R}_{2}$ are consistent with UAT. However, for $UAT^{R}_{2}$, its parameter $\mathbf{b}^{\prime}_{2,2}$ is influenced by the input, while other weight parameters, such as $\mathbf{W}^{\prime}_{2,1}$, can be disregarded. These parameters are only affected by other parameters, so once the network training is complete, they are essentially fixed. This allows us to focus primarily on the dynamic parameters. In equation $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, except for the initial term, $\mathbf{b}^{\prime}_{2,2}$ also follows the same mathematical form as UAT. This can be interpreted as dynamically adjusting the bias term $\mathbf{b}^{\prime}_{2,2}$ via UAT based on the input. In conclusion, we have demonstrated that the mathematical forms of one-layer and two-layer residual networks are consistent with the UAT framework and the two-layer residual network is the DUAT function.
Next, we will use mathematical induction to prove that the mathematical form of a multi-layer residual network is also a DUAT function. Assume that the overall mathematical form of the first $i$ layers of the residual network aligns with UAT. Our goal is to show that the overall mathematical form of the $i+1$-th layer remains consistent with UAT.
@@
For clarity, we make the following definitions: since the overall mathematical form of the first $i$ layers is consistent with UAT, we can write it as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, where the first term is explicitly written, and the remainder is denoted as $UAT^{R}_{i}$, with $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$. Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
@@
First, consider $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$. Substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, we obtain Eq. 26. Setting $\mathbf{W}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}$ and $\mathbf{b}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$. Since the overall mathematical form of $UAT^{R}_{i}$ is consistent with the $i$-layer UAT and we have shown that the UAT form is preserved when multiplied by a matrix, we have thus demonstrated that the overall mathematical form of the first part remains consistent with UAT.